# Robust Design and Evaluation of

### v5

@@ Appendix A Proofs of Results in the Main Text
A.5 Proof of Proposition 4.1
In order to prove Proposition 4.1, we consider the infeasible case in which the bounding functions $\underline{\delta}(\cdot;\eta),\overline{\delta}(\cdot;\eta)$ are known as a stepping stone. We prove the result for the upper bound since the argument is identical for the lower bound,
@@
where $\Delta_{n}=\left\{\tilde{\delta}\colon(1-D_{i})\underline{\delta}(X_{i};\eta)\leq\tilde{\delta}_{i}\leq(1-D_{i})\overline{\delta}(X_{i};\eta)\mbox{ for }i=1,\ldots,n/3\right\}$ and $\underline{\delta}(X_{i};\eta)=(\underline{\Gamma}-1)\mu_{1}(X_{i})$, $\overline{\delta}(X_{i};\eta)=(\overline{\Gamma}-1)\mu_{1}(X_{i})$. As shorthand, define $\underline{\delta}^{\prime}(D_{i},X_{i};\eta)=(1-D_{i})\underline{\delta}(X_{i};\eta)$ and $\overline{\delta}^{\prime}(D_{i},X_{i};\eta)=(1-D_{i})\overline{\delta}(X_{i};\eta)$.
Lemma A.1.
Define $\mathcal{U}$ to be the set of monotone, non-decreasing functions $u(\cdot)\colon\mathbb{R}\rightarrow[0,1]$ and $\Delta^{M}:=\left\{\delta^{\prime}(d,x)=\underline{\delta}^{\prime}(d,x;\eta)+(\overline{\delta}^{\prime}(d,x;\eta)-\underline{\delta}^{\prime}(d,x;\eta))u(\beta_{0}(x))\mbox{ for }u(\cdot)\in\mathcal{U}\right\}$. Let $\Delta^{M}_{n}=\{(\delta^{\prime}(D_{1},X_{1}),\ldots,\delta^{\prime}(D_{n/3},X_{n/3})\colon\delta\in\Delta^{M}\}$. Then, $\overline{\mbox{perf}}_{+}(s;\beta,\Delta):=\sup_{\tilde{\delta}\in\Delta^{M}}\mbox{perf}_{+}(s;\beta,\tilde{\delta})$ and
@@
s.t. $1\displaystyle 0\leq\tilde{U}_{i}\leq\tilde{V}\mbox{ for }i=1,\ldots n/3,\mbox{}0\leq\tilde{V},\mbox{ }\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}=$ 1 .
@@
Next, notice that the constraint $-\textbf{1}^{\prime}P+\lambda\hat{d}\geq\hat{c}$ must be tight at the optimum. Plugging in the previous expression for $P_{i}$ and the expressions for $\hat{c},\hat{d}$, this implies that $\lambda$ satisfies
@@
    $-\mathbb{E}_{n}[(\beta_{0,i}-\lambda)(\overline{\delta}^{\prime}_{i}-\underline{\delta}^{\prime}_{i})]=\mathbb{E}_{n}[\left(\beta_{0,i}-\lambda\right)\left(\phi_{\mu,i}(\hat{\eta})+\underline{\delta}^{\prime}_{i}\right)]\implies\lambda=\frac{\mathbb{E}_{n}[\beta_{0,i}\left(\phi_{\mu,i}(\hat{\eta})+\overline{\delta}^{\prime}_{i}\right)]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\overline{\delta}^{\prime}_{i}]}.$
@@
The constraint $\langle\textbf{1},\tilde{P}\rangle_{P(\cdot)}+\lambda d\geq c$ must be tight at the optimum. Plugging in the previous expression for $\tilde{P}(\cdot)$, this implies that $\lambda$ satisfies
@@
$\left\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})-\overline{\mbox{perf}}_{+}(s;\beta,\Delta)\right\|$ $=\left\|\sup_{\tilde{\delta}\in\Delta_{n}^{M}}\widehat{\mbox{perf}}_{+}(s;\beta,\tilde{\delta})-\sup_{\tilde{\delta}\in\Delta^{M}}\mbox{perf}_{+}(s;\beta,\tilde{\delta})\right\|$
$=\left\|\sup_{\tilde{\delta}\in\Delta^{M}}\widehat{\mbox{perf}}_{+}(s;\beta,\tilde{\delta})-\sup_{\tilde{\delta}\in\Delta^{M}}\mbox{perf}{+}(s;\beta,\tilde{\delta})\right\|$
$\leq\sup_{\tilde{\delta}\in\Delta^{M}}\left\|\widehat{\mbox{perf}}_{+}(s;\beta,\tilde{\delta})-\mbox{perf}_{+}(s;\beta,\tilde{\delta})\right\|,$
where the first equality uses Lemma A.1. For any $\tilde{\delta}\in\Delta^{M}$, we have that
$\widehat{\mbox{perf}}_{+}(s;\beta,\tilde{\delta})-\mbox{perf}_{+}(s;\beta,\tilde{\delta})=$
@@
$\frac{\mathbb{E}_{n}[(1)]}{\mathbb{E}_{n}[(2)]}-\frac{\mathbb{E}[(3)]}{\mathbb{E}[(4)]}=\mathbb{E}_{n}[(2)]^{-1}\left\{\mathbb{E}_{n}[(1)]-\mathbb{E}[(3)]-\frac{\mathbb{E}[(3)]}{\mathbb{E}[(4)]}(\mathbb{E}_{n}[(2)]-\mathbb{E}[(4)])\right\},$
@@
$\mathbb{E}_{n}[(1)]-\mathbb{E}[(3)]$ $=\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]-\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)+\beta_{0,i}\tilde{\delta}_{i}]$
@@
$\mathbb{E}_{n}[(2)]-\mathbb{E}[(4)]$ $=\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]-\mathbb{E}[\phi_{\mu,i}(\eta)+\tilde{\delta}_{i}]$
@@
$\mathbb{E}_{n}[(2)]=\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]\geq\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}]$
$\mathbb{E}[(3)]=\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)+\beta_{0,i}\tilde{\delta}_{i}]\leq\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)+\beta_{0,i}(1-D_{i})\overline{\delta}_{i}]$
$\mathbb{E}[(4)]=\mathbb{E}[\mu_{1}(X_{i})+\tilde{\delta}_{i}]\geq\mathbb{E}[\mu_{1}(X_{i})+(1-D_{i})\underline{\delta}_{i}].$
Therefore, there exists $C_{1}>0$ such that $\mathbb{E}_{n}[(2)]>C_{1}$ for all $n$ under the assumption of bounded nuisance parameter estimates. There also exists constants $C_{2}<\infty$, $C_{3}>0$ such that $\mathbb{E}[(3)]<C_{2}$ and $\mathbb{E}[(4)]>C_{3}$. Putting this together, we therefore have
@@
$\sup_{U\in\mathcal{U}}\left|(n/3)^{-1}\sum_{i}\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(\beta_{0,i})-\mathbb{E}[\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(\beta_{0,i})]\right|.$
Define $f(w,x,y,z)=w(y-x)z$ and =$\mathcal{F}=\{f_{U}\}_{U\in\mathcal{U}}$ to be the class of functions $f_{U}\colon(d,\underline{\delta},\overline{\delta},\beta)\rightarrow(1-d)(\overline{\delta}-\underline{\delta})\beta u(\beta)$. Observe $f$ is a contraction over its final argument on $[0,1]$. We can then rewrite (c) as
$\sup_{f_{U}\in\mathcal{F}}\left|(n/3)^{-1}\sum_{i}f_{U}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})-\mathbb{E}[f_{\tilde{\delta}}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})]\right|.$
@@
$\sup_{f_{U}\in\mathcal{F}}\left|(n/3)^{-1}\sum_{i}f_{U}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})-\mathbb{E}[f_{U}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})]\right|\leq R_{n}(\mathcal{F})+\sqrt{\frac{2\log(1/\delta)}{n/3}},$
where $R_{n}(\mathcal{F})$ is the Rademacher complexity of $\mathcal{F}$. Now we relate $R_{n}(\mathcal{F})$ to $R_{n}(\mathcal{U})$. For any fixed $(d_{1},\underline{\delta}_{1},\overline{\delta}_{i},\beta_{0,1}),\ldots,(d_{n/3},\underline{\delta}_{n/3},\overline{\delta}_{n/3},\beta_{0,n/3})$, observe that
$\mathbb{E}_{\epsilon}[\sup_{U\in\mathcal{U}}\left|\sum_{i=1}^{n/3}\epsilon_{i}f_{U}(d_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})\right|]$ $=\mathbb{E}_{\epsilon}[\sup_{U\in\mathcal{U}}\left|\sum_{i=1}^{n/3}\epsilon_{i}\beta_{0,i}(1-d_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(\beta_{0,i})\right|]$
$\leq 2(\overline{\Gamma}-\underline{\Gamma})\mathbb{E}_{\epsilon}[\sup_{U\in\mathcal{U}}\left|\sum_{i=1}^{n/3}\epsilon_{i}U(\beta_{0,i})\right|]$
where we used that $(1-d_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})\leq(\overline{\Gamma}-\underline{\Gamma})$ and the Ledoux-Talagrand contraction inequality (e.g., Eq. (5.61) in Wainwright_textbook). Dividing by $n/3$ and averaging over the observations yields $R_{n}(\mathcal{F})\leq 2(\overline{\Gamma}-\underline{\Gamma})R_{n}(\mathcal{U})$. Finally, we can bound the Rademacher complexity of $\mathcal{U}$ using Dudley’s entropy integral (e.g., Theorem 5.22 in Wainwright_textbook) as
@@
[…]
$\sup_{U\in\mathcal{U}}\left\|(\mathbb{E}_{n}^{k}-\mathbb{E})[\beta_{0,i}(\overline{\delta}_{i}^{\prime}-\underline{\delta}_{i}^{\prime})U(\beta_{0,i})]\right\|\leq\frac{C}{\sqrt{n/3}}+\sqrt{\frac{2\log(1/\delta)}{2n/3}}$
Now return to Proposition 4.1 in which the researcher specifies nonparametric outcome bounds. We prove a more general lemma about the estimator for the upper bound on positive class performance with estimated bounding functions $\hat{\overline{\delta}}(\cdot),\hat{\underline{\delta}}(\cdot)$ and then show that it implies Proposition 4.1. Suppose we solve the following maximization problem in each fold
@@
Lemma A.2 analyzed the convergence rate of the second term. It is therefore sufficient to bound the first term.
Lemma A.3.
Assume the same conditions as Lemma A.2 and $P(\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\hat{\underline{\delta}}_{i})>c)=1$ for some $c>0$. Then,
$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|\lesssim\sqrt{\frac{1}{n/3}\sum_{i=1}^{n/3}(\widehat{\underline{\delta}}_{i}-\underline{\delta}_{i})^{2}}+\sqrt{\frac{1}{n/3}\sum_{i=1}^{n/3}(\widehat{\overline{\delta}}_{i}-\overline{\delta}_{i})^{2}},$
@@
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n}):=\max_{0\leq U\leq 1}$ $\frac{\sum_{i=1}^{n/3}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\underline{\delta}_{i}+\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}{\sum_{i=1}^{n_{k}}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}+(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}$
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n}):=\max_{0\leq U\leq 1}$ $\frac{\sum_{i=1}^{n/3}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\hat{\underline{\delta}}_{i}+\beta_{0,i}(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}{\sum_{i=1}^{n/3}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\hat{\underline{\delta}}_{i}+(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}.$
@@
$\max_{0\leq U\leq 1}\left\|\frac{\sum_{i=1}^{n/3}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\hat{\underline{\delta}}_{i}+\beta_{0,i}(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}{\sum_{i=1}^{n/3}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\hat{\underline{\delta}}_{i}+(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}-\frac{\sum_{i=1}^{n/3}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\underline{\delta}_{i}+\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}{\sum_{i=1}^{n/3}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}+(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}\right\|=$
$\max_{0\leq U\leq 1}\|\frac{\mathbb{E}_{n}[(1)]}{\mathbb{E}_{n}[(2)]}-\frac{\mathbb{E}_{n}[(3)]}{\mathbb{E}_{n}[(4)]}\|=\max_{0\leq U\leq 1}\mathbb{E}_{n}[(2)]^{-1}\left\{\underbrace{\left(\mathbb{E}_{n}[(1)]-\mathbb{E}_{n}[(3)]\right)}_{(a)}-\frac{\mathbb{E}_{n}[(3)]}{\mathbb{E}_{n}[(4)]}\underbrace{\left(\mathbb{E}_{n}[(2)]-\mathbb{E}_{n}[(4)]\right)}_{(b)}\right\}.$

### v6

@@ Appendix B Proofs of Results in the Main Text
B.5 Proof of Proposition 4.1
To prove Proposition 4.1, we first state and prove various auxiliary lemmas. Let $\mathcal{O}_{k}$ denote the observations in the $k$-th fold and $\mathcal{O}_{-k}$ denote the observations not in the $k$-th fold.
Lemma B.1.
Proof.
$\mathbb{E}\left[\beta(X_{i})\frac{(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))(\hat{\pi}_{1}(X_{i})-\pi_{1}(X_{i}))}{\hat{\pi}_{1}(X_{i})}\right]\overset{(2)}{\leq}\\$
Lemma B.2 (Lemma 2 in KennedyEtAl(2020)-SharpInstruments).
Lemma B.3.
Proof.
[…]
Proof.
This result follows directly from the stated conditions after some algebra. Suppressing dependence on the subscript k-k to ease notation, observe that we can rewrite
[…]
Proof.
This result follows directly from the stated conditions after some simple algebra. We omit the dependence on $X_{i}$ and k-k. Observe that we can rewrite
[…]
$\|\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\|\overset{(5)}{=}$
$\|\hat{\pi}_{0}(\hat{\mu}_{1}-\mu_{1})-\mu_{1}(\pi_{0}-\hat{\pi}_{0})\|$
The result then follows. ∎
Lemma B.6.
Proof.
For ease of notation, we omit the dependence on $X_{i}$ and the subscript k-k. The proof follows an analogous argument to Lemma B.1. Observe that
where
where (c) is $O_{\mathbb{P}}(\|\hat{\pi}_{1,-k}-\pi_{1}\|/\sqrt{n}+\|\hat{\mu}_{1,-k}+\mu_{1}\|/\sqrt{n}+R_{1,n}^{k}/\sqrt{n})$ by Lemma B.3 and Lemma B.4, and (d) is also $O_{\mathbb{P}}(\|\hat{\pi}_{1,-k}-\pi_{1}\|/\sqrt{n}+\|\hat{\mu}_{1,-k}+\mu_{1}\|/\sqrt{n}+R_{1,n}^{k}/\sqrt{n})$ Lemma B.3 and Lemma B.5. The second term $\mathbb{E}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}\mid\mathcal{O}_{k}]|$ is bounded by
where (e) is $O_{\mathbb{P}}(R_{1,n}^{k})$ by Lemma B.1 and (f) is $O_{\mathbb{P}}(R_{1,n}^{k})$ by Lemma B.6. Putting this together, we have shown the first claim
if $R_{1,n}=o_{\mathbb{P}}(1/\sqrt{n})$. By the central limit theorem,
$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)\xrightarrow{d}N\left(0,Cov\left(\begin{pmatrix}\overline{\mbox{perf}}_{i}\\ \underline{\mbox{perf}}_{i}\end{pmatrix}\right)\right)$,
from which the result follows. $\Box$
B.6 Proof of Proposition 4.2
To prove Proposition 4.2, we first consider the case in which $\underline{\delta}(\cdot;\eta),\overline{\delta}(\cdot;\eta)$ are known as a stepping stone. We prove the result for the upper bound since the argument is identical for the lower bound. To simplify notation, we assume the data are split into two equal sized folds.
@@
where $\Delta_{n}=\left\{\tilde{\delta}\colon(1-D_{i})\underline{\delta}(X_{i};\eta)\leq\tilde{\delta}_{i}\leq(1-D_{i})\overline{\delta}(X_{i};\eta)\mbox{ for }i=1,\ldots,n/2\right\}$ and $\underline{\delta}(X_{i};\eta)=(\underline{\Gamma}-1)\mu_{1}(X_{i})$, $\overline{\delta}(X_{i};\eta)=(\overline{\Gamma}-1)\mu_{1}(X_{i})$. As shorthand, define $\underline{\delta}^{\prime}(D_{i},X_{i};\eta)=(1-D_{i})\underline{\delta}(X_{i};\eta)$ and $\overline{\delta}^{\prime}(D_{i},X_{i};\eta)=(1-D_{i})\overline{\delta}(X_{i};\eta)$.
Lemma B.7.
[…]
@@
s.t. $0\leq\tilde{U}_{i}\leq\tilde{V}\mbox{ for }i=1,\ldots n/2,\mbox{ }0\leq\tilde{V},\mbox{ }\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}=1.$
@@
Next, notice that the constraint $\textbf{1}^{\prime}P+\lambda\hat{d}\geq\hat{c}$ must be tight at the optimum. Plugging in the previous expression for $P_{i}$ and the expressions for $\hat{c},\hat{d}$, this implies that $\lambda$ satisfies
@@
    $\mathbb{E}_{n}[(\beta_{0,i}-\lambda)(\overline{\delta}^{\prime}_{i}-\underline{\delta}^{\prime}_{i})]=\mathbb{E}_{n}[\left(\beta_{0,i}-\lambda\right)\left(\phi_{\mu,i}(\hat{\eta})+\underline{\delta}^{\prime}_{i}\right)]\implies\lambda=\frac{\mathbb{E}_{n}[\beta_{0,i}\left(\phi_{\mu,i}(\hat{\eta})+\overline{\delta}^{\prime}_{i}\right)]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\overline{\delta}^{\prime}_{i}]}.$
@@
[…]
@@
[…]
where the first equality uses Lemma B.7. For any $\tilde{\delta}\in\Delta^{M}$, we have that
[…]
@@
$\frac{\mathbb{E}_{n}[(\text{\#1})]}{\mathbb{E}_{n}[(\text{\#2})]}-\frac{\mathbb{E}[(\text{\#3})]}{\mathbb{E}[(\text{\#4})]}=\mathbb{E}_{n}[(\text{\#2})]^{-1}\left\{\mathbb{E}_{n}[(\text{\#1})]-\mathbb{E}[(\text{\#3})]-\frac{\mathbb{E}[(\text{\#3})]}{\mathbb{E}[(\text{\#4})]}(\mathbb{E}_{n}[(\text{\#2})]-\mathbb{E}[(\text{\#4})])\right\},$
@@
$\mathbb{E}_{n}[(\text{\#1})]-\mathbb{E}[(\text{\#3})]$ $=\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]-\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)+\beta_{0,i}\tilde{\delta}_{i}]$
@@
$\mathbb{E}_{n}[(\text{\#2})]-\mathbb{E}[(\text{\#4})]$ $=\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]-\mathbb{E}[\phi_{\mu,i}(\eta)+\tilde{\delta}_{i}]$
@@
$\mathbb{E}_{n}[(\text{\#2})]=\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]\geq\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}]$
$\mathbb{E}[(\text{\#3})]=\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)+\beta_{0,i}\tilde{\delta}_{i}]\leq\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)+\beta_{0,i}(1-D_{i})\overline{\delta}_{i}]$
$\mathbb{E}[(\text{\#4})]=\mathbb{E}[\mu_{1}(X_{i})+\tilde{\delta}_{i}]\geq\mathbb{E}[\mu_{1}(X_{i})+(1-D_{i})\underline{\delta}_{i}].$
Therefore, there exists $C_{1}>0$ such that $\mathbb{E}_{n}[(\text{\#2})]>C_{1}$ for all $n$ under the assumption of bounded nuisance parameter estimates. There also exists constants $C_{2}<\infty$, $C_{3}>0$ such that $\mathbb{E}[(\text{\#3})]<C_{2}$ and $\mathbb{E}[(\text{\#4})]>C_{3}$. Putting this together, we therefore have
@@
$\sup_{U\in\mathcal{U}}\left|(n/2)^{-1}\sum_{i}\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(\beta_{0,i})-\mathbb{E}[\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(\beta_{0,i})]\right|.$
Define f(w,x,y,z)=w(yx)zf(w,x,y,z)=w(y-x)z and $\mathcal{F}=\{f_{U}\}_{U\in\mathcal{U}}$ to be the class of functions $f_{U}\colon(d,\underline{\delta},\overline{\delta},\beta)\rightarrow(1-d)(\overline{\delta}-\underline{\delta})\beta u(\beta)$. Observe $f$ is a contraction over its final argument on $[0,1]$. We can then rewrite (c) as
$\sup_{f_{U}\in\mathcal{F}}\left|(n/2)^{-1}\sum_{i}f_{U}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})-\mathbb{E}[f_{\tilde{\delta}}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})]\right|.$
@@
$\sup_{f_{U}\in\mathcal{F}}\left|(n/2)^{-1}\sum_{i}f_{U}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})-\mathbb{E}[f_{U}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})]\right|\leq R_{n}(\mathcal{F})+\sqrt{\frac{2\log(1/\delta)}{n/2}},$
where $R_{n}(\mathcal{F})$ is the Rademacher complexity of $\mathcal{F}$. Now we relate $R_{n}(\mathcal{F})$ to $R_{n}(\mathcal{U})$. For any fixed $(d_{1},\underline{\delta}_{1},\overline{\delta}_{i},\beta_{0,1}),\ldots,(d_{n/2},\underline{\delta}_{n/2},\overline{\delta}_{n/2},\beta_{0,n/2})$, observe that
$\mathbb{E}_{\epsilon}[\sup_{U\in\mathcal{U}}\left|\sum_{i=1}^{n/2}\epsilon_{i}f_{U}(d_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})\right|]$ $=\mathbb{E}_{\epsilon}[\sup_{U\in\mathcal{U}}\left|\sum_{i=1}^{n/2}\epsilon_{i}\beta_{0,i}(1-d_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(\beta_{0,i})\right|]$
$\leq 2(\overline{\Gamma}-\underline{\Gamma})\mathbb{E}_{\epsilon}[\sup_{U\in\mathcal{U}}\left|\sum_{i=1}^{n/2}\epsilon_{i}U(\beta_{0,i})\right|]$
where we used that $(1-d_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})\leq(\overline{\Gamma}-\underline{\Gamma})$ and the Ledoux-Talagrand contraction inequality (e.g., Eq. (5.61) in Wainwright_textbook). Dividing by $n/2$ and averaging over the observations yields $R_{n}(\mathcal{F})\leq 2(\overline{\Gamma}-\underline{\Gamma})R_{n}(\mathcal{U})$. Finally, we can bound the Rademacher complexity of $\mathcal{U}$ using Dudley’s entropy integral (e.g., Theorem 5.22 in Wainwright_textbook) as
$R_{n}(\mathcal{U})\leq\frac{C}{\sqrt{n/2}}\int_{0}^{1}\sqrt{\log(N(\xi,\mathcal{U},\|\cdot\|_{\mathbb{P}_{n}}))}d\xi\leq\frac{C}{\sqrt{n/2}}\int_{0}^{1}\sqrt{\log(N_{[]}(2\xi,\mathcal{U},\|\cdot\|_{\mathbb{P}_{n}}))}d\xi,$
for some constant $C$, where $N(\xi,\mathcal{U},\|\cdot\|_{\mathbb{P}_{n}})$ is the covering number and $N_{[]}(2\xi,\mathcal{U},\|\cdot\|_{p_{n}})$ is the bracketing number. But, Theorem 2.7.5 of VDVWellner establishes that the bracketing entropy $\log(N_{[]}(\xi,\mathcal{U},\|\cdot\|_{\mathbb{P}_{n}})$ of the class of monotone non-decreasing functions is bounded by $(1/\xi)\log(1/\xi)$, and so $\int_{0}^{1}\sqrt{\log(N_{[]}(\xi,\mathcal{U},\|\cdot\|_{\mathbb{P}_{n}}))}d\xi=\sqrt{2\pi}$. It follows that, for any $\delta>0$,
$\sup_{U\in\mathcal{U}}\left\|(\mathbb{E}_{n}^{k}-\mathbb{E})[\beta_{0,i}(\overline{\delta}_{i}^{\prime}-\underline{\delta}_{i}^{\prime})U(\beta_{0,i})]\right\|\leq\frac{C}{\sqrt{n/2}}+\sqrt{\frac{2\log(1/\delta)}{2n/2}}$
@@
Lemma B.8 analyzed the convergence rate of the second term. It is therefore sufficient to bound the first term.
Lemma B.9.
Assume the same conditions as Lemma B.8. Then,
$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|\lesssim\sqrt{\frac{1}{n/2}\sum_{i=1}^{n/2}(\widehat{\underline{\delta}}_{i}-\underline{\delta}_{i})^{2}}+\sqrt{\frac{1}{n/2}\sum_{i=1}^{n/2}(\widehat{\overline{\delta}}_{i}-\overline{\delta}_{i})^{2}},$
@@
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n}):=\max_{0\leq U\leq 1}$ $\frac{\sum_{i=1}^{n/2}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\underline{\delta}_{i}+\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}{\sum_{i=1}^{n_{k}}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}+(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}$
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n}):=\max_{0\leq U\leq 1}$ $\frac{\sum_{i=1}^{n/2}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\hat{\underline{\delta}}_{i}+\beta_{0,i}(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}{\sum_{i=1}^{n/2}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\hat{\underline{\delta}}_{i}+(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}.$
@@
$\max_{0\leq U\leq 1}\left\|\frac{\sum_{i=1}^{n/2}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\hat{\underline{\delta}}_{i}+\beta_{0,i}(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}{\sum_{i=1}^{n/2}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\hat{\underline{\delta}}_{i}+(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}-\frac{\sum_{i=1}^{n/2}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\underline{\delta}_{i}+\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}{\sum_{i=1}^{n/2}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}+(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}\right\|=$
$\max_{0\leq U\leq 1}\|\frac{\mathbb{E}_{n}[(\text{\#1})]}{\mathbb{E}_{n}[(\text{\#2})]}-\frac{\mathbb{E}_{n}[(\text{\#3})]}{\mathbb{E}_{n}[(\text{\#4})]}\|=\max_{0\leq U\leq 1}\mathbb{E}_{n}[(\text{\#2})]^{-1}\left\{\underbrace{\left(\mathbb{E}_{n}[(\text{\#1})]-\mathbb{E}_{n}[(\text{\#3})]\right)}_{(a)}-\frac{\mathbb{E}_{n}[(\text{\#3})]}{\mathbb{E}_{n}[(\text{\#4})]}\underbrace{\left(\mathbb{E}_{n}[(\text{\#2})]-\mathbb{E}_{n}[(\text{\#4})]\right)}_{(b)}\right\}.$